If I connect a large number of 75.0 microFarads capacitors in parallel across a 260.0 Volt battery, roughly how many capacitors do I need to store 180.0 Joules of energy?
Answer #1
     Given
    voltage V = 260 V
    energy U = 180 J
     energy stored in the capacitors
            U = (1/2) CV2   
             180    = 0.5 * C* 2602
                C = 5325μF
   So we need 5325μF capacitance for 180 J ofenergy
    we have each capacitor has75μF  
            number of capacitors C = 5325/75 = 71
   so 71 capacitors are connected in parallel for 180 J of energy